OpenStudy (anonymous):

I'm having trouble solving a differential equation using Laplace Transforms. y'''' - 2y'' + y = 0 After applying the Laplace Transforms, I get: L{y} = 4(s^2 + 1) / (s^2 - 1)^2 But I don't know how to perform the inverse Laplace Transform on it. Please advise.

OpenStudy (anonymous):

u got the formula for the laplace transform for your 2nd and 3rd derivative?

OpenStudy (anonymous):

to solve ode with laplace transform u will need initial conditions, y(0), y'(0) and y''(0)

OpenStudy (anonymous):

Yeah, y(0) = 0, y'(0) = 4, y''(0) = 0, y'''(0) = 8. The thing is, I already did those, I just don't know how to do the inverse Laplace. I've looked on wolfram and I saw the answer, but I want to know how to get to it.

OpenStudy (anonymous):

u know how to do partial fractions? apply partial fraction and see what u can get?

OpenStudy (anonymous):

factor in denominator\[(ax^{2}+bx+c)^{k}=\frac{A _{1}x+B _{1}}{ax^{2}+bx+c}+\frac{A _{2}x+B _{2}}{(ax^{2}+bx+c)^{2}}+...+\frac{A _{k}x+B _{k}}{(ax^{2}+bx+c)^{k}}\]

OpenStudy (anonymous):

I'm not sure partial fractions works with a squared term as the denominator, though.

OpenStudy (anonymous):

u can factor out your denominator into linear terms b4 applying partial fraction,\[(x^{2}-1)^{2}=(x-1)^{2}(x+1)^{2}\]so now you are solving\[L[y]=\frac{A}{x-1}+\frac{Bs+C}{(x-1)^{2}}+\frac{D}{x+1}+\frac{Es+F}{(x+1)^{2}}\]

OpenStudy (anonymous):

i get C = F = 2, other unknowns are 0, so\[L[y]=\frac{2}{(x-1)^{2}}+\frac{2}{(x+1)^{2}}\]

OpenStudy (anonymous):

Oh wow, I see. Okay. Thank you so much! :D

OpenStudy (anonymous):

Wait, can you explain how you got C = F = 2? I'm getting B + C = 2 and F - E = 2, and also -A + C + D + F = 4.

OpenStudy (anonymous):

sry my laptop shutdown for no reason suddenly

OpenStudy (amistre64):

wow, a y'''' , someone must really love you

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

im so sorry man, its A B C D only

OpenStudy (anonymous):

._.\[\frac{A}{x-1}+\frac{B}{(x-1)^{2}}+\frac{C}{x+1}+\frac{D}{(x+1)^{2}}\]so B and D =2, A=C=0. very sorry for my mistake.

OpenStudy (anonymous):

Yup, I got to that step. And @amistre64 at least it's not any higher :P

OpenStudy (amistre64):

Im sure that there is a pattern that can be spotted after y''' y^n = s^n Y(s) - (y(0)+sy'(0)+s^2y''(0)+s^3y'''(0)...s^(n-1)y^(n-1)(0)) but i would have to verify that to be sure

OpenStudy (anonymous):

Yeah, that's a theorem. But I've already gone past that step, it's just the partial fraction decomposition that was bothering me.

OpenStudy (anonymous):

collect the coefficients for the term with the same power and compare it with 4(s^2 +1)

OpenStudy (anonymous):

But with that method, I'm only ending up with 2 solutions. I should have 4, since it's a fourth order differential equation. I think the answers are e^t, te^t, e^-t, and te^-t. Also, if the Laplace was \[L\left[ y \right] = 2/(x+1)^2 + 2/(x-1)2\], wouldn't those answers come out to be 2te^t and 2te^-t instead of te^t and te^-t?

OpenStudy (turingtest):

for reference purposes here is the solution, though the wolf chose not to use laplace http://www.wolframalpha.com/input/?i=solve%20y''''-2y''%2By%3D0%2Cy(0)%20%3D%200%2C%20y'(0)%20%3D%204%2C%20y''(0)%20%3D%200%2C%20y'''(0)%20%3D%208&t=crmtb01